# Hierarchical Transformer Preconditioning for Interactive Physics Simulation

### 13 May 2026 v1

@@ Abstract.
The key training contribution is a cosine-Hutchinson probe objective that learns the action of MAMA on convergence-critical spectral subspaces, optimizing angular alignment of $MA\mathbf{z}$ with $\mathbf{z}$ rather than forcing eigenvalue clusters to a prescribed location. This removes unnecessary spectral-placement constraints from SAI-style objectives and improves conditioning on irregular spectra. Because both inference and apply are dense, dependency-free tensor programs, the full solve loop is captured as a single CUDA Graph.
@@
copyright: nonejournalyear: 2026doi: 10.1145/nnnnnnn.nnnnnnnjournal: TOGjournalvolume: 45journalnumber: 6article: 1publicationmonth: 12ccs: Computing methodologies Physical simulationccs: Computing methodologies Neural networks
@@ 1. Introduction
Real-time simulation of fluids, soft bodies, and coupled multiphysics often reduces every time step to a large sparse SPD solve attacked with preconditioned conjugate gradient (PCG) (Saad, 2003). The preconditioner $M\!\approx\!A^{-1}$ is the dominant design variable. Classical choices — Jacobi, incomplete Cholesky/LU (IC/ILU), sparse approximate inverses (SPAI), and algebraic multigrid (AMG) (Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987) — excel when the matrix is reused across many solves so a heavy setup amortizes. In an interactive setting AA changes every frame, so any preconditioner whose setup approaches the per-frame budget cannot amortize; IC and AMG suffer worst (Naumov et al., 2015; Liu et al., 2016; Yamazaki et al., 2020).
Machine learning offers an alternative: train a neural preconditioner once and amortize the cost across frames. Graph neural networks have been the dominant choice (Li et al., 2023; Häusner et al., 2023; Chen, 2024; Yang et al., 2025) because a sparse matrix has a natural graph interpretation, but the choice forces two limitations that become acute in real-time simulation. (i) GNN preconditioners typically inherit the sparsity of AA, so non-adjacent interactions are either dropped or only reached through repeated graph convolutions that oversmooth local detail (Wu et al., 2019; Trifonov et al., 2025). (ii) The dominant operations at apply time — triangular solves on a learned IC, or two SpMVs on a learned SPAI — are bandwidth-bound and either serialize on data hazards or scatter through irregular memory.
@@
The ideal preconditioner is $A^{-1}$, which is dense even when AA is sparse, so the question is what structure makes $A^{-1}$ cheap to store and apply. In modern simulation codes the degrees of freedom are already laid out along a space-filling curve (Morton, Hilbert) or a bandwidth-reducing permutation, because the surrounding pipeline (spatial hashing, BVH, neighbor search) wants that ordering (Karras, 2012; Teschner et al., 2003; Ihmsen et al., 2011). Under such orderings the nonzeros of AA cluster near the diagonal and the corresponding off-diagonal blocks of $A^{-1}$ describe long-range interactions whose numerical rank decays with separation. Hierarchical matrices ($\mathcal{H}$-matrices) (Hackbusch, 1999; Hackbusch and Khoromskij, 2002; Börm et al., 2003) exploit exactly this — a recursive block partition with dense diagonal blocks and low-rank off-diagonal blocks — and the fast multipole method (Greengard and Rokhlin, 1987) factors the same prior for translation-invariant kernels. An $\mathcal{H}$-matrix preconditioner is architecturally an FMM-shaped operator on the assembled system, and that is the prior we hand the network.
@@
We train self-supervised with a cosine-similarity Hutchinson probe loss on the preconditioned image $MA\mathbf{z}$. The conceptual move over the Frobenius/SAI loss used by (Yang et al., 2025) is to replace a distance between vectors ($\|\tfrac{1}{\|A\|}AM\mathbf{z}-\mathbf{z}\|_{2}$) with a distance between subspaces (the angle between the lines spanned by $\mathbf{z}$ and $MA\mathbf{z}$). This is invariant under positive rescaling of $M$, exactly matching PCG, which cares about the direction MAMA sends each probe and not the magnitude. The network is then free to cluster the spectrum of MAMA wherever angular alignment is easiest, with no implicit demand that the cluster live near any particular eigenvalue (Prop. 6.1; spectral evidence in Fig. 2). Empirically, this is not a minor objective tweak: with architecture and training split fixed, the loss swap alone accounts for a major quality gap in Table 6.
@@
In short: [c1] a neural preconditioner that bakes the $\mathcal{H}$-matrix block partition in as a structural prior with a fixed coarse rank on every off-diagonal tile, keeping total work at $\Theta(N)$ (intro paragraph c1, §4.2); [c2] highway buffers that route information between blocks without growing the budget (§4.3); [c3] a cosine-similarity Hutchinson loss that minimizes a distance between subspaces rather than between vectors, exactly matching PCG’s scale-invariance and freeing the network to put the spectrum of MAMA wherever convergence is easiest (intro paragraph c3, Prop. 6.1 in §6.1); [c4] an allocation-free apply with no triangular solves, no data-dependent branching, no per-kernel CPU dispatch — the whole PCG inner loop captured as one CUDA Graph (intro paragraph c4, §5). Table 1 summarizes the per-iteration access pattern and throughput limit of each preconditioner family at the one step where every fixed-pattern alternative loses time on modern GPUs.
@@ 2. Related Work
Jacobi, incomplete Cholesky/LU, sparse approximate inverses (SPAI), and algebraic multigrid (Saad, 2003; Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987; Naumov et al., 2015) are mature and remain default choices when the system is reused across many solves so a careful setup amortizes. The known GPU costs of triangular solves (Liu et al., 2016; Yamazaki et al., 2020) and of V-cycle communication (Naumov et al., 2015) bite hardest in the regime we target (AA changing every frame at interactive rates), which motivates the amortized neural alternative pursued here.
@@ 3. Background
so CG only sees AA through how small a polynomial fixed to 11 at the origin can be made on $\sigma(A)$. The practical consequence we lean on is that when $\sigma(MA)$ collapses to $c$ tight clusters away from zero, CG converges in $\sim\!c$ iterations regardless of where on the real line those clusters sit. The preconditioning objective is therefore “cluster $\sigma(MA)$,” not “make MAMA close to II,” with cluster location essentially free — the property our cosine loss exploits.
@@
The training loss uses Hutchinson-style probes $\mathbf{z}$ with $\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]\!=\!I$ (Hutchinson, 1989), but not as a trace estimator. The property we use is spectral whiteness: isotropy gives $\mathbb{E}[(\mathbf{u}_{i}^{\top}\mathbf{z})^{2}]\!=\!1$ for every eigenvector $\mathbf{u}_{i}$ of MAMA, so a scalar built from $(\mathbf{z},MA\mathbf{z})$ weighs every eigenmode equally rather than favoring smooth or oscillatory ones — the property that makes the cosine loss of c3 a faithful global indicator of preconditioner quality.
@@ 4. Method
$A\!\in\!\mathbb{R}^{N\times N}$ is the assembled sparse SPD system, indexed along a space-filling-curve or bandwidth-reducing ordering. The NN indices are partitioned into K=N/LK\!=\!N/L contiguous leaves of size $L$, inducing an $\mathcal{H}$-matrix partition of $K$ diagonal blocks and $M_{\mathcal{H}}$ unique weakly-admissible off-diagonal tiles (tile $m$ spans $S_{m}\!\times\!S_{m}$ leaves). The network is a four-stage pipeline: a physics-aware encoder, a diagonal attention stack (full per-node resolution $L$), an off-diagonal attention stack (coarse token count $L_{s}\!\ll\!L$, the same for every tile), and linear decoder heads. The whole network dispatches exactly two attention kernels regardless of NN — one batched over leaves, one batched over tiles — which, together with the static partition, is what later lets the entire PCG inner loop run inside a single CUDA Graph. We use $L\!=\!128,L_{s}\!=\!32,d\!=\!128,n_{l}\!=\!3$ in every reported result.
@@ 4.1. Encoder
A two-layer MLP lifts whichever per-node features the simulator already exposes (position, velocity, density, pressure, material parameters) plus the geometric and coupling features $(\Delta\mathbf{x}_{ij},A_{ij})$ for every nonzero of AA to width-$d$ node embeddings; $n_{\mathrm{gcn}}\!=\!2$ residual graph-convolutional layers (Kipf and Welling, 2017) mix in neighborhood information using AA as the message-passing weight. This is the only stage that observes individual graph edges; everything downstream operates on the block layout of the partition, decoupling cost from $\mathrm{nnz}(A)$. Because $M$ is neural rather than algebraic, the network can look at $(\rho_{i},\Delta\mathbf{x}_{ij})$ directly — physical structure a fixed analytical preconditioner cannot see (full feature list in Supplementary LABEL:SX-supp:architecture).
@@ 4.2. Diagonal and off-diagonal attention stacks
Off-diagonal stack. For each off-diagonal tile $m$ of size $S_{m}L\!\times\!S_{m}L$, node embeddings on its row- and column-strips are pooled to a fixed coarse token count $L_{s}\!=\!L/p_{\mathrm{off}}\!\ll\!L$ regardless of $S_{m}$. $n_{l}$ transformer layers over these $L_{s}$ tokens emit a factor pair $U_{m},V_{m}\!\in\!\mathbb{R}^{L_{s}\times L_{s}}$; the tile’s coarse representation is $B_{m}\!=\!U_{m}V_{m}^{\top}$. Because the token count is the same on every tile, the implied rank fraction $L_{s}/(S_{m}L)$ shrinks automatically with separation — exactly the $\mathcal{H}$-matrix prior that distant blocks need less rank. A rank-fraction audit on real frames confirms the architecture-provided $L_{s}/(SL)$ stays above what truncated SVD needs at $\varepsilon\!\in\!\{10^{-3},10^{-6},10^{-9}\}$ in every distance class (Fig. 3; assembled-$M$ visualization in Supplementary, Fig. LABEL:SX-supp-fig:matrices). With $S_{m}$ doubling geometrically, $M_{\mathcal{H}}\!=\!O(K)$ unique off-diagonal tiles, and a tile-batched attention of fixed shape, the total work over the off-diagonal stack stays linear in NN.
@@ 4.3. Highway connections
$4.$ Decoder, geometric picture, and complexity
@@ 5. Preconditioner Application
Every kernel in the inner loop has fixed launch shape, no host-side allocation, and no data-dependent control flow, so the entire iteration — preconditioner apply, SpMV, SAXPYs, dot-product reductions — records into one CUDA Graph that subsequent iterations replay with a single cudaGraphLaunch. CPU dispatch between kernels disappears, which in a real-time engine usually is the floor on per-frame latency. The graph-capturable property falls out of the design rather than being engineered: IC/DILU triangular solves serialize through data-dependent wavefronts and AMG V-cycles need level-by-level barriers, so neither is single-graph capturable in the same way. Among learned alternatives, neural SPAI (Yang et al., 2025) is graph-capturable but applies $\mathbf{y}\!=\!G^{\top}(G\mathbf{r})$ as two SpMVs against a sparse $G$ whose nonzero pattern follows AA, so its apply remains bandwidth-bound on irregular gathers; in contrast, our factorized approximate-inverse preconditioner apply reduces to dense block GEMMs on the partition and is compute-bound on Tensor Cores.
@@ 6.1. Cosine Hutchinson Probe Loss
We train the network self-supervised. Given a batch of probe vectors $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ drawn from the spectrally-balanced distribution described below, we compute $A\mathbf{Z}$ via SpMV (treated as a fixed input, with no gradient through AA), apply the preconditioner to obtain $MA\mathbf{Z}$, and minimize the angle between $MA\mathbf{Z}$ and $\mathbf{Z}$ as flattened tensors:
@@
The conceptual upgrade $\mathcal{L}_{\cos}$ makes over Frobenius- and SAI-style objectives is to replace a distance between vectors with a distance between subspaces. Frobenius-type losses — including the SAI loss $\bigl\|\tfrac{1}{\|A\|}AM-I\bigr\|_{F}^{2}$ used by (Yang et al., 2025) — penalize the pointwise Euclidean deviation of the specific vector $\tfrac{1}{\|A\|}AM\mathbf{z}$ from the specific vector $\mathbf{z}$. That distance is sensitive to the magnitude of $AM\mathbf{z}$, so it implicitly demands that AMAM act as identity at a particular absolute scale — in the SAI case, at scale $\|A\|$. $\mathcal{L}_{\cos}$, by contrast, sees only the direction of $MA\mathbf{Z}$. Two preconditioners that send a probe to the same one-dimensional subspace incur the same loss, no matter how they scale the vector inside that subspace. Geometrically, $\mathcal{L}_{\cos}$ is a distance on the projective space $\mathbb{P}(\mathbb{R}^{NK_{z}})$ rather than a Euclidean distance on $\mathbb{R}^{NK_{z}}$:
@@
Let $M:\mathbb{R}^{N}\!\to\!\mathbb{R}^{N}$ be a linear preconditioner, AA a fixed SPD matrix, and $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ a probe matrix with $\mathbf{Z}\!\neq\!0$ and $MA\mathbf{Z}\!\neq\!0$. Write $\widehat{\mathbf{Z}}\!=\!\operatorname{vec}(\mathbf{Z})$ and $\widehat{\mathbf{Y}}\!=\!\operatorname{vec}(MA\mathbf{Z})$, both elements of $\mathbb{R}^{NK_{z}}$.
@@
The proposition matches the PCG geometry one-to-one. PCG’s convergence depends only on the relative spread of the eigenvalues of MAMA, not on their absolute location (Sec. 3); accordingly, the correct space to optimize $M$ over is the projective space $M/\mathbb{R}_{>0}$, and the natural loss on that space is a distance between the subspaces $\mathbb{R}\widehat{\mathbf{Z}}$ and $\mathbb{R}\widehat{\mathbf{Y}}$ they span, exactly what $\mathcal{L}_{\cos}$ provides. Frobenius- and SAI-style losses live on the wrong space — they pin down the absolute scale of MAMA even though PCG does not care — and as a side effect implicitly demand that the eigenvalues of MAMA cluster near a chosen value ($\|A\|$ in the SAI case), wasting capacity on a constraint with no algorithmic payoff. By dropping that constraint, $\mathcal{L}_{\cos}$ frees the network to cluster the spectrum wherever the current preconditioner makes angular alignment easiest — a behavior we observe directly in §7.3 (Figs. 42) and ablate against SAI in §7.3.

### 14 May 2026 v2

@@ Abstract.
The key training contribution is a cosine-Hutchinson probe objective that learns the action of $MA$ on convergence-critical spectral subspaces, optimizing angular alignment of $MA\mathbf{z}$ with $\mathbf{z}$ rather than forcing eigenvalue clusters to a prescribed location. This removes unnecessary spectral-placement constraints from SAI-style objectives and improves conditioning on irregular spectra. Because both inference and apply are dense, dependency-free tensor programs, the full solve loop is captured as a single CUDA Graph.
@@
copyright: noneccs: Computing methodologies Physical simulationccs: Computing methodologies Neural networks
@@ 1. Introduction
Real-time simulation of fluids, soft bodies, and coupled multiphysics often reduces every time step to a large sparse SPD solve attacked with preconditioned conjugate gradient (PCG) (Saad, 2003). The preconditioner $M\!\approx\!A^{-1}$ is the dominant design variable. Classical choices — Jacobi, incomplete Cholesky/LU (IC/ILU), sparse approximate inverses (SPAI), and algebraic multigrid (AMG) (Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987) — excel when the matrix is reused across many solves so a heavy setup amortizes. In an interactive setting $A$ changes every frame, so any preconditioner whose setup approaches the per-frame budget cannot amortize; IC and AMG suffer worst (Naumov et al., 2015; Liu et al., 2016; Yamazaki et al., 2020).
Machine learning offers an alternative: train a neural preconditioner once and amortize the cost across frames. Graph neural networks have been the dominant choice (Li et al., 2023; Häusner et al., 2023; Chen, 2024; Yang et al., 2025) because a sparse matrix has a natural graph interpretation, but the choice forces two limitations that become acute in real-time simulation. (i) GNN preconditioners typically inherit the sparsity of $A$, so non-adjacent interactions are either dropped or only reached through repeated graph convolutions that oversmooth local detail (Wu et al., 2019; Trifonov et al., 2025). (ii) The dominant operations at apply time — triangular solves on a learned IC, or two SpMVs on a learned SPAI — are bandwidth-bound and either serialize on data hazards or scatter through irregular memory.
@@
The ideal preconditioner is $A^{-1}$, which is dense even when $A$ is sparse, so the question is what structure makes $A^{-1}$ cheap to store and apply. In modern simulation codes the degrees of freedom are already laid out along a space-filling curve (Morton, Hilbert) or a bandwidth-reducing permutation, because the surrounding pipeline (spatial hashing, BVH, neighbor search) wants that ordering (Karras, 2012; Teschner et al., 2003; Ihmsen et al., 2011). Under such orderings the nonzeros of $A$ cluster near the diagonal and the corresponding off-diagonal blocks of $A^{-1}$ describe long-range interactions whose numerical rank decays with separation. Hierarchical matrices ($\mathcal{H}$-matrices) (Hackbusch, 1999; Hackbusch and Khoromskij, 2002; Börm et al., 2003) exploit exactly this — a recursive block partition with dense diagonal blocks and low-rank off-diagonal blocks — and the fast multipole method (Greengard and Rokhlin, 1987) factors the same prior for translation-invariant kernels. An $\mathcal{H}$-matrix preconditioner is architecturally an FMM-shaped operator on the assembled system, and that is the prior we hand the network.
@@
We train self-supervised with a cosine-similarity Hutchinson probe loss on the preconditioned image $MA\mathbf{z}$. The conceptual move over the Frobenius/SAI loss used by (Yang et al., 2025) is to replace a distance between vectors ($\|\tfrac{1}{\|A\|}AM\mathbf{z}-\mathbf{z}\|_{2}$) with a distance between subspaces (the angle between the lines spanned by $\mathbf{z}$ and $MA\mathbf{z}$). This is invariant under positive rescaling of $M$, exactly matching PCG, which cares about the direction $MA$ sends each probe and not the magnitude. The network is then free to cluster the spectrum of $MA$ wherever angular alignment is easiest, with no implicit demand that the cluster live near any particular eigenvalue (Prop. 6.1; spectral evidence in Fig. 2). Empirically, this is not a minor objective tweak: with architecture and training split fixed, the loss swap alone accounts for a major quality gap in Table 6.
@@
In short: [c1] a neural preconditioner that bakes the $\mathcal{H}$-matrix block partition in as a structural prior with a fixed coarse rank on every off-diagonal tile, keeping total work at $\Theta(N)$ (intro paragraph c1, §4.2); [c2] highway buffers that route information between blocks without growing the budget (§4.3); [c3] a cosine-similarity Hutchinson loss that minimizes a distance between subspaces rather than between vectors, exactly matching PCG’s scale-invariance and freeing the network to put the spectrum of $MA$ wherever convergence is easiest (intro paragraph c3, Prop. 6.1 in §6.1); [c4] an allocation-free apply with no triangular solves, no data-dependent branching, no per-kernel CPU dispatch — the whole PCG inner loop captured as one CUDA Graph (intro paragraph c4, §5). Table 1 summarizes the per-iteration access pattern and throughput limit of each preconditioner family at the one step where every fixed-pattern alternative loses time on modern GPUs.
@@ 2. Related Work
Jacobi, incomplete Cholesky/LU, sparse approximate inverses (SPAI), and algebraic multigrid (Saad, 2003; Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987; Naumov et al., 2015) are mature and remain default choices when the system is reused across many solves so a careful setup amortizes. The known GPU costs of triangular solves (Liu et al., 2016; Yamazaki et al., 2020) and of V-cycle communication (Naumov et al., 2015) bite hardest in the regime we target ($A$ changing every frame at interactive rates), which motivates the amortized neural alternative pursued here.
@@ 3. Background
so CG only sees $A$ through how small a polynomial fixed to $1$ at the origin can be made on $\sigma(A)$. The practical consequence we lean on is that when $\sigma(MA)$ collapses to $c$ tight clusters away from zero, CG converges in $\sim\!c$ iterations regardless of where on the real line those clusters sit. The preconditioning objective is therefore “cluster $\sigma(MA)$,” not “make $MA$ close to II,” with cluster location essentially free — the property our cosine loss exploits.
@@
The training loss uses Hutchinson-style probes $\mathbf{z}$ with $\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]\!=\!I$ (Hutchinson, 1989), but not as a trace estimator. The property we use is spectral whiteness: isotropy gives $\mathbb{E}[(\mathbf{u}_{i}^{\top}\mathbf{z})^{2}]\!=\!1$ for every eigenvector $\mathbf{u}_{i}$ of $MA$, so a scalar built from $(\mathbf{z},MA\mathbf{z})$ weighs every eigenmode equally rather than favoring smooth or oscillatory ones — the property that makes the cosine loss of c3 a faithful global indicator of preconditioner quality.
@@ 4. Method
$A\!\in\!\mathbb{R}^{N\times N}$ is the assembled sparse SPD system, indexed along a space-filling-curve or bandwidth-reducing ordering. The $N$ indices are partitioned into K=N/LK\!=\!N/L contiguous leaves of size $L$, inducing an $\mathcal{H}$-matrix partition of $K$ diagonal blocks and $M_{\mathcal{H}}$ unique weakly-admissible off-diagonal tiles (tile $m$ spans $S_{m}\!\times\!S_{m}$ leaves). The network is a four-stage pipeline: a physics-aware encoder, a diagonal attention stack (full per-node resolution $L$), an off-diagonal attention stack (coarse token count $L_{s}\!\ll\!L$, the same for every tile), and linear decoder heads. The whole network dispatches exactly two attention kernels regardless of $N$ — one batched over leaves, one batched over tiles — which, together with the static partition, is what later lets the entire PCG inner loop run inside a single CUDA Graph. We use $L\!=\!128,L_{s}\!=\!32,d\!=\!128,n_{l}\!=\!3$ in every reported result.
@@ 4.1. Encoder
A two-layer MLP lifts whichever per-node features the simulator already exposes (position, velocity, density, pressure, material parameters) plus the geometric and coupling features $(\Delta\mathbf{x}_{ij},A_{ij})$ for every nonzero of $A$ to width-$d$ node embeddings; $n_{\mathrm{gcn}}\!=\!2$ residual graph-convolutional layers (Kipf and Welling, 2017) mix in neighborhood information using $A$ as the message-passing weight. This is the only stage that observes individual graph edges; everything downstream operates on the block layout of the partition, decoupling cost from $\mathrm{nnz}(A)$. Because $M$ is neural rather than algebraic, the network can look at $(\rho_{i},\Delta\mathbf{x}_{ij})$ directly — physical structure a fixed analytical preconditioner cannot see (full feature list in Supplementary LABEL:SX-supp:architecture).
@@ 4.2. Diagonal and off-diagonal attention stacks
Off-diagonal stack. For each off-diagonal tile $m$ of size $S_{m}L\!\times\!S_{m}L$, node embeddings on its row- and column-strips are pooled to a fixed coarse token count $L_{s}\!=\!L/p_{\mathrm{off}}\!\ll\!L$ regardless of $S_{m}$. $n_{l}$ transformer layers over these $L_{s}$ tokens emit a factor pair $U_{m},V_{m}\!\in\!\mathbb{R}^{L_{s}\times L_{s}}$; the tile’s coarse representation is $B_{m}\!=\!U_{m}V_{m}^{\top}$. Because the token count is the same on every tile, the implied rank fraction $L_{s}/(S_{m}L)$ shrinks automatically with separation — exactly the $\mathcal{H}$-matrix prior that distant blocks need less rank. A rank-fraction audit on real frames confirms the architecture-provided $L_{s}/(SL)$ stays above what truncated SVD needs at $\varepsilon\!\in\!\{10^{-3},10^{-6},10^{-9}\}$ in every distance class (Fig. 3; assembled-$M$ visualization in Supplementary, Fig. LABEL:SX-supp-fig:matrices). With $S_{m}$ doubling geometrically, $M_{\mathcal{H}}\!=\!O(K)$ unique off-diagonal tiles, and a tile-batched attention of fixed shape, the total work over the off-diagonal stack stays linear in $N$.
@@ 4.3. Highway connections
4.4. Decoder, geometric picture, and complexity
@@ 5. Preconditioner Application
Every kernel in the inner loop has fixed launch shape, no host-side allocation, and no data-dependent control flow, so the entire iteration — preconditioner apply, SpMV, SAXPYs, dot-product reductions — records into one CUDA Graph that subsequent iterations replay with a single cudaGraphLaunch. CPU dispatch between kernels disappears, which in a real-time engine usually is the floor on per-frame latency. The graph-capturable property falls out of the design rather than being engineered: IC/DILU triangular solves serialize through data-dependent wavefronts and AMG V-cycles need level-by-level barriers, so neither is single-graph capturable in the same way. Among learned alternatives, neural SPAI (Yang et al., 2025) is graph-capturable but applies $\mathbf{y}\!=\!G^{\top}(G\mathbf{r})$ as two SpMVs against a sparse $G$ whose nonzero pattern follows $A$, so its apply remains bandwidth-bound on irregular gathers; in contrast, our factorized approximate-inverse preconditioner apply reduces to dense block GEMMs on the partition and is compute-bound on Tensor Cores.
@@ 6.1. Cosine Hutchinson Probe Loss
We train the network self-supervised. Given a batch of probe vectors $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ drawn from the spectrally-balanced distribution described below, we compute $A\mathbf{Z}$ via SpMV (treated as a fixed input, with no gradient through $A$), apply the preconditioner to obtain $MA\mathbf{Z}$, and minimize the angle between $MA\mathbf{Z}$ and $\mathbf{Z}$ as flattened tensors:
@@
The conceptual upgrade $\mathcal{L}_{\cos}$ makes over Frobenius- and SAI-style objectives is to replace a distance between vectors with a distance between subspaces. Frobenius-type losses — including the SAI loss $\bigl\|\tfrac{1}{\|A\|}AM-I\bigr\|_{F}^{2}$ used by (Yang et al., 2025) — penalize the pointwise Euclidean deviation of the specific vector $\tfrac{1}{\|A\|}AM\mathbf{z}$ from the specific vector $\mathbf{z}$. That distance is sensitive to the magnitude of $AM\mathbf{z}$, so it implicitly demands that $AM$ act as identity at a particular absolute scale — in the SAI case, at scale $\|A\|$. $\mathcal{L}_{\cos}$, by contrast, sees only the direction of $MA\mathbf{Z}$. Two preconditioners that send a probe to the same one-dimensional subspace incur the same loss, no matter how they scale the vector inside that subspace. Geometrically, $\mathcal{L}_{\cos}$ is a distance on the projective space $\mathbb{P}(\mathbb{R}^{NK_{z}})$ rather than a Euclidean distance on $\mathbb{R}^{NK_{z}}$:
@@
Let $M:\mathbb{R}^{N}\!\to\!\mathbb{R}^{N}$ be a linear preconditioner, $A$ a fixed SPD matrix, and $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ a probe matrix with $\mathbf{Z}\!\neq\!0$ and $MA\mathbf{Z}\!\neq\!0$. Write $\widehat{\mathbf{Z}}\!=\!\operatorname{vec}(\mathbf{Z})$ and $\widehat{\mathbf{Y}}\!=\!\operatorname{vec}(MA\mathbf{Z})$, both elements of $\mathbb{R}^{NK_{z}}$.
@@
The proposition matches the PCG geometry one-to-one. PCG’s convergence depends only on the relative spread of the eigenvalues of $MA$, not on their absolute location (Sec. 3); accordingly, the correct space to optimize $M$ over is the projective space $M/\mathbb{R}_{>0}$, and the natural loss on that space is a distance between the subspaces $\mathbb{R}\widehat{\mathbf{Z}}$ and $\mathbb{R}\widehat{\mathbf{Y}}$ they span, exactly what $\mathcal{L}_{\cos}$ provides. Frobenius- and SAI-style losses live on the wrong space — they pin down the absolute scale of $MA$ even though PCG does not care — and as a side effect implicitly demand that the eigenvalues of $MA$ cluster near a chosen value ($\|A\|$ in the SAI case), wasting capacity on a constraint with no algorithmic payoff. By dropping that constraint, $\mathcal{L}_{\cos}$ frees the network to cluster the spectrum wherever the current preconditioner makes angular alignment easiest — a behavior we observe directly in §7.3 (Figs. 42) and ablate against SAI in §7.3.